# FreeInsert: Disentangled Text-Guided Object Insertion in 3D Gaussian Scene without Spatial Priors

### v3

@@ 1. Introduction
Text-driven 3D generation (Poole et al., 2022; Zhou et al., 2024b; Bokhovkin et al., 2025) and editing (Xu et al., 2024; Zhou et al., 2024a; Barda et al., 2025) are gaining traction for enabling the intuitive customization of digital content with a few words. Despite recent advances (Dong and Wang, 2023; Haque et al., 2023; Mirzaei et al., 2025; Khalid et al., 2024; Wang et al., 2024a; Zhuang et al., 2023; Song et al., 2023) that have made significant progress in editing the geometry and appearance of scene components, flexibly inserting new objects into the scene remains challenging due to difficulties in precise placement and seamless integration.
@@
Recent 3D editing methods leverage diffusion models by first performing text-guided 2D edits on single (Shahbazi et al., 2024; Chen et al., 2024a) or multi-view images (Haque et al., 2023; Zhuang et al., 2023), then lifting them to 3D. Relying solely on textual descriptions for object insertion often leads to insertion failure and suboptimal results due to misinterpretation of the text (Zhuang et al., 2024, 2023), as shown in Figure 3, exemplified by methods like Instruct-NeRF2NeRF (Haque et al., 2023) and GaussCtrl (Wu et al., 2024). Some methods introduce attention mechanisms to capture the spatial relationship between the inserted object and the scene. However, these methods still struggle with accurately determining the inserted object’s pose and scale (Zhuang et al., 2023; Sun et al., 2024; Zhou et al., 2024a). To address this limitation, other methods leverage user-provided 2D masks (Chen et al., 2024a; Shahbazi et al., 2024) or 3D bounding boxes (Zhuang et al., 2024; Shum et al., 2024) as strong constraints to achieve more controllable and concise insertion. Nevertheless, they often demand specialized expertise (Zhuang et al., 2024) and considerable manual effort, limiting their practical usability. In addition, they still face challenges with inaccurate depth estimation (Chen et al., 2024a) and inconsistent 3D multi-view reconstruction due to the modality gap between 2D and 3D. We summarize the above discussion as shown in Table 1.
@@
Inspired by this, achieving flexible and high-quality object insertion into 3D scenes without manual supervision remains underexplored. The advent of large-scale models (Achiam et al., 2023; Deitke et al., 2024; Xiao et al., 2024; Liu et al., 2024; Grattafiori et al., 2024), which have acquired human commonsense knowledge, has made unsupervised learning increasingly promising. In this paper, we propose FreeInsert, a method that leverages foundation models (MLLMs (Achiam et al., 2023; Deitke et al., 2024), LGM (Tang et al., 2024) and Diffusion model (Rombach et al., 2022) ) to assist object insertion in 3D scenes without relying on any spatial priors. Our method removes the need for spatial priors by inferring object insertion directly from high-level textual cues (e.g., “Add [object] to/on [target]”) as Figure 1 shows. We argue that the insertion process can essentially be viewed as first generating a object, followed by estimating the transformation, which defines the inserted object’s degrees of freedom (pose and scale) relative to the scene.
@@
Specifically, we disentangle the object insertion process into object generation and its parameterized degrees of freedom (DoF) estimation, both guided by textual descriptions with foundation models. We first obtain a text instruction from the user and parse it into structured semantics (e.g., object type, spatial relation, attachment region) using an MLLM-based (Achiam et al., 2023) object insertion parser. This enables precise, controllable object insertion that aligns with the user’s intent. We then employ a 3D-consistent reconstruction model (Tang et al., 2024) to obtain an initial Gaussian-based object model, which is coarsely inserted into the scene guided by the visual and spatial reasoning capabilities of MLLMs(Achiam et al., 2023; Deitke et al., 2024). This step inherently circumvents the inconsistency issues associated with 2D editing-based methods. While the feed-forward procedure provides a lightweight 3D layout, it often suffers from suboptimal placement and imperfect geometry. To address these issues, we propose a two-stage refinement. First, the Hierarchical Spatial-Aware Refinement stage optimizes the object’s DoF via spatially-aware score distillation sampling (SSDS) (Chen et al., 2024b) from pretrained diffusion model (Rombach et al., 2022). This stage leverages MLLM-derived reasoning results to align the object’s DoF with both local and global spatial semantics, enhancing more precise and controllable placement. These reasoning results also help the model handle rare spatial composition e.g., “Add a pair of sunglasses on the forehead”, thereby improving robustness. In the final appearance refinement stage, we fine-tune a pretrained diffusion model on multi-view renderings of the optimized inserted object and its corresponding inserted-object image, and use it to enhance the object’s appearance. By disentangling placement from object generation, our method enables flexible semantic control over insertion while preserving object quality and ensuring coherent, plausible 3D scenes.
@@
No Required Manual Supervision 3D View Consistency Support Image-Prompts Instruct-N2N(Haque et al., 2023) $\times$ $\times$ GaussCtrl (Wu et al., 2024) $\times$ GaussianEditor (Chen et al., 2024a) $\times$ $\times$ $\times$ TIP-Editor (Zhuang et al., 2024) $\times$ $\times$ FreeInsert
@@ 2. Related Works
Text-Guided 3D Scene Editing. Text-driven 3D scene editing has seen rapid progress, thanks to the rise of diffusion models (Rombach et al., 2022; Ho et al., 2020). Most methods (Kim et al., 2023; Koo et al., 2024; Haque et al., 2023; Mirzaei et al., 2025; Park et al., 2024; Dong and Wang, 2023; Wang et al., 2024b) focus on modifying existing content, either globally or locally. Local editing requires precise localization to avoid affecting unrelated regions, which remains challenging. While some works use implicit cues from models like InstructPix2Pix (Brooks et al., 2023) or ControlNet (Zhang et al., 2023), others incorporate explicit constraints such as segmentation masks (Wang et al., 2024a) or cross-attention maps (Zhuang et al., 2023). However, these methods struggle with 3D object insertion, which demands reasoning about semantically appropriate yet physically unoccupied regions for placement. In this work, we primarily focus on the task of object insertion in 3D scenes.
@@
Object Insertion in 3D Scene. In contrast to modifying existing scene content, object insertion remains underexplored. MVInpainter (Cao et al., [n. d.]) leverages segmentation to identify support regions like table surfaces, but struggles with fine-grained insertions on objects or humans. GaussianEditor (Chen et al., 2024a) and InseRF (Shahbazi et al., 2024) insert objects using 3D reconstruct models, yet still require user-provided 2D masks and suffer from depth-related localization issues. FocalDreamer (Li et al., 2024) attaches parts to base shapes but depends on user-specified 3D parameters (e.g., rotation, translation, scale), and lacks generalization to complex scenes. Other methods (Zhuang et al., 2024; Shum et al., 2024) guide object generation via diffusion using 3D bounding boxes, which imposes a burden on users. In this work, we aim for unsupervised and broadly applicable 3D object insertion, removing the need for manual annotations or spatial priors.
@@
Large Language Models in 3D Generation and Editing. LLMs, like GPT (Achiam et al., 2023) and Llama (Grattafiori et al., 2024) series, have exhibited outstanding efficacy in many text-related tasks. Zhou et al. (2024b) and Zhou et al. (2025) utilize LLMs to provide coarse compositional spatial information from textual descriptions to construct the 3D scene. The multi-modal variants of LLMs (Deitke et al., 2024; Achiam et al., 2023) incorporate images and are additionally trained on image-text pairs, showing impressive results for visual captioning and vision question-answering (VQA). Notably, Molmo (Deitke et al., 2024) can perform pixel-level localization mainly because it was trained with richly annotated image data. This capability is crucial for robust spatial grounding in vision-language tasks. GG-Editor (Xu et al., 2024) first exploits GPT-4V (Achiam et al., 2023) to better understand both the textual and 3D visual inputs and then infer reasonable local regions for 3D editing. However, it primarily targets object editing and its preliminary use of MLLM struggles to ensure spatial precision. In this work, we leverage the text reasoning capabilities and spatial relationship understanding of GPT-4 (Achiam et al., 2023) and Molmo (Deitke et al., 2024), supplemented by a basic detection model (Xiao et al., 2024), to eliminate the reliance on manually provided priors in object insertion.
@@ 3.2. MLLM-based Object Insertion Parser
A key challenge in unsupervised object insertion is converting high-level user intent into structured, fine-grained guidance. To address this, we introduce an MLLM-based Object Insertion Parser (MLLM-OIP) that utilizes the MLLM’s spatial understanding capability to parse the instruction $\mathcal{T}$ into semantically prompts, providing essential guidance for the subsequent object insertion. Specifically, we provide a prompt template $\mathcal{T}_{\textit{parser}}$ and a sampled scene image $\mathcal{I}_{\textit{S}}$ as input to the multimodal LLM $\mathcal{M}_{\textit{MLLM}}$ (Achiam et al., 2023) to obtain structured outputs. The prompts generation process is formalized as :
@@ 3.3. Initialization via Large Models
Object from Prompts. To avoid 3D inconsistency, we first use a text-to-image (T2I) (Rombach et al., 2022) model to synthesize a Text-generated image $\mathcal{I}_{\textit{O}}$ of the object from the object description prompt $\mathcal{T}_{\textit{O}}$. The synthesized image is then used to recover the 3D geometry $\mathcal{G}_{O}$ via LGM (Tang et al., 2024), a single-view reconstruction model that achieve a trade-off between reconstruction quality and efficiency. Other lightweight 3D reconstruction methods (Long et al., 2024; Hong et al., 2023) can also be adopted. In addition, $\mathcal{I}_{\textit{O}}$ can be directly specified by the user, allowing for more precise control over the object’s appearance.
Scene’s Attachment Region Detection. Intuitively, an object’s placement is influenced by the attachment region of the scene and the degrees of freedom within that region. Driven by this, we extract an attachment region $\mathcal{G}_{\textit{AR}}$ and its associated degrees of freedom $\textit{DoF}_{\textit{AR}}$($\textit{s}_{\textit{AR}}$, $\textit{r}_{\textit{AR}}$, $\textit{t}_{\textit{AR}}$) from the 3D scene based on the Attachment Region Prompt $\mathcal{T}_{\textit{AR}}$. This region serves as a crucial spatial reference, guiding initializing the inserted object $\textit{DoF}_{\textit{O}}$. Specifically, we employ an open-vocabulary detection model Florence2 (Xiao et al., 2024) to localize 2D bounding boxes across smapled views from the scene with camera poses $\mathcal{C}_{\textit{cam}}$, guided by $\mathcal{T}_{\textit{AR}}$. For each view, the detected box is converted into a binary mask $\mathcal{I}_{\textit{BAR}}$, representing the candidate attachment region. The 3D attachment area is parameterized by the degrees of freedom of a initial 3D bounding box $\mathcal{B}_{\textit{init}}$. We optimize the attachment by computing cross-entropy between the projected transformed bounding box and the detected attachment region mask $\mathcal{B}_{\text{AR}}$ across all camera views. Thus, $\textit{DoF}_{\textit{AR}}$ is calculated as follows:
@@
To initialize the $\textit{t}_{\textit{O}}$, we use strong pixel-level semantic spatial localization capability of Molmo (Deitke et al., 2024) to predict a set of 2D object centers $\{c_{O}^{(v)}\}_{v\in\mathcal{V}}$ across multiple scene views with the Local Target $\mathcal{T}_{\textit{LT}}$, using prompt like “Point the position to add $\mathcal{T}_{\textit{LT}}$¿”. Let $\hat{\mathcal{G}}_{O}^{(t)}$ denote the object geometry after applying the transformation $(\textit{s}_{\textit{O}}$, $\textit{r}_{\textit{O}}$, $t$ ), where $t$ is a optimized parameter. For each view $v$, we project the transformed object and compute the 2D centroid of its projection. The $\textit{t}_{\textit{O}}$ is obtained by optimizing $t$ to minimize the discrepancy between the projected and the predicted centroids:
@@
Here, $\pi_{v}(\cdot)$ is the camera projection function for view $v$, and Centroid ($\cdot$) computes the 2D projected center of the object. To ensure physical plausibility during object insertion, we introduce the collision loss (Zhou et al., 2025) $\mathcal{L}_{\text{coll}}$, which penalizes interpenetration between the object centriod $O_{c}$ and the scene attachment region $\mathcal{G}_{\text{AR}}$.
@@ 3.4. Hierarchical Spatial Aware Refinement
The initial $\textit{DoF}_{\textit{O}}$ from $\mathcal{M}_{\textit{MLLM}}$ often lack spatial accuracy, hindering seamless scene integration. Base on that, we then optimize the $\textit{DoF}_{\textit{O}}$ using SSDS Loss (Chen et al., 2024b), refining the object’s placement in the scene. The loss is defined as:
@@
Attention-based Localization. We found that due to the bias in the T2I model’s training data, SSDS Loss exhibits limitations when handling rare spatial relationships. These inherent limitations restrict the effectiveness of prompt instructions. To enable stronger spatial conditioning, we adopt attention-based localization loss (Zhuang et al., 2024), enforcing tighter regional constraints as follows:

### v4

@@ 1. Introduction
Text-driven 3D generation (poole2022dreamfusiontextto3dusing2d; zhou2024gala3d; bokhovkin2024scenefactor) and editing (xu2024gg; zhou2024edit3d; barda2024instant3dit) are gaining traction for enabling the intuitive customization of digital content with a few words. Despite recent advances (dong2023vica; haque2023instruct; mirzaei2025watch; khalid2024latenteditor; wang2024gaussianeditor; zhuang2023dreameditor; song2023blending) that have made significant progress in editing the geometry and appearance of scene components, flexibly inserting new objects into the scene remains challenging due to difficulties in precise placement and seamless integration.
@@
Recent 3D editing methods leverage diffusion models by first performing text-guided 2D edits on single (shahbazi2024inserf; chen2024gaussianeditor) or multi-view images (haque2023instruct; zhuang2023dreameditor), then lifting them to 3D. Relying solely on textual descriptions for object insertion often leads to insertion failure and suboptimal results due to misinterpretation of the text (zhuang2024tip; zhuang2023dreameditor), as shown in Figure 3, exemplified by methods like Instruct-NeRF2NeRF (haque2023instruct) and GaussCtrl (wu2024gaussctrl). Some methods introduce attention mechanisms to capture the spatial relationship between the inserted object and the scene. However, these methods still struggle with accurately determining the inserted object’s pose and scale (zhuang2023dreameditor; sun2024gseditpro; zhou2024edit3d). To address this limitation, other methods leverage user-provided 2D masks (chen2024gaussianeditor; shahbazi2024inserf) or 3D bounding boxes (zhuang2024tip; shum2024language) as strong constraints to achieve more controllable and concise insertion. Nevertheless, they often demand specialized expertise (zhuang2024tip) and considerable manual effort, limiting their practical usability. In addition, they still face challenges with inaccurate depth estimation (chen2024gaussianeditor) and inconsistent 3D multi-view reconstruction due to the modality gap between 2D and 3D. We summarize the above discussion as shown in Table 1.
@@
Inspired by this, achieving flexible and high-quality object insertion into 3D scenes without manual supervision remains underexplored. The advent of large-scale models (achiam2023gpt; deitke2024molmo; xiao2024florence; liu2024deepseek; grattafiori2024llama), which have acquired human commonsense knowledge, has made unsupervised learning increasingly promising. In this paper, we propose FreeInsert, a method that leverages foundation models (MLLMs (achiam2023gpt; deitke2024molmo), LGM (tang2024lgm) and Diffusion model (rombach2022high) ) to assist object insertion in 3D scenes without relying on any spatial priors. Our method removes the need for spatial priors by inferring object insertion directly from high-level textual cues (e.g., “Add [object] to/on [target]”) as Figure 1 shows. We argue that the insertion process can essentially be viewed as first generating a object, followed by estimating the transformation, which defines the inserted object’s degrees of freedom (pose and scale) relative to the scene.
@@
Specifically, we disentangle the object insertion process into object generation and its parameterized degrees of freedom (DoF) estimation, both guided by textual descriptions with foundation models. We first obtain a text instruction from the user and parse it into structured semantics (e.g., object type, spatial relation, attachment region) using an MLLM-based (achiam2023gpt) object insertion parser. This enables precise, controllable object insertion that aligns with the user’s intent. We then employ a 3D-consistent reconstruction model (tang2024lgm) to obtain an initial Gaussian-based object model, which is coarsely inserted into the scene guided by the visual and spatial reasoning capabilities of MLLMs(achiam2023gpt; deitke2024molmo). This step inherently circumvents the inconsistency issues associated with 2D editing-based methods. While the feed-forward procedure provides a lightweight 3D layout, it often suffers from suboptimal placement and imperfect geometry. To address these issues, we propose a two-stage refinement. First, the Hierarchical Spatial-Aware Refinement stage optimizes the object’s DoF via spatially-aware score distillation sampling (SSDS) (chen2024comboverse) from pretrained diffusion model (rombach2022high). This stage leverages MLLM-derived reasoning results to align the object’s DoF with both local and global spatial semantics, enhancing more precise and controllable placement. These reasoning results also help the model handle rare spatial composition e.g., “Add a pair of sunglasses on the forehead”, thereby improving robustness. In the final appearance refinement stage, we fine-tune a pretrained diffusion model on multi-view renderings of the optimized inserted object and its corresponding inserted-object image, and use it to enhance the object’s appearance. By disentangling placement from object generation, our method enables flexible semantic control over insertion while preserving object quality and ensuring coherent, plausible 3D scenes.
@@
No Required Manual Supervision 3D View Consistency Support Image-Prompts Instruct-N2N(haque2023instruct) $\times$ $\times$ GaussCtrl (wu2024gaussctrl) $\times$ GaussianEditor (chen2024gaussianeditor) $\times$ $\times$ $\times$ TIP-Editor (zhuang2024tip) $\times$ $\times$ FreeInsert
@@ 2. Related Works
Text-Guided 3D Scene Editing. Text-driven 3D scene editing has seen rapid progress, thanks to the rise of diffusion models (rombach2022high; ho2020denoising). Most methods (kim2023collaborative; koo2024posterior; haque2023instruct; mirzaei2025watch; park2024ednerf; dong2023vica; wang2024innerf360) focus on modifying existing content, either globally or locally. Local editing requires precise localization to avoid affecting unrelated regions, which remains challenging. While some works use implicit cues from models like InstructPix2Pix (brooks2023instructpix2pix) or ControlNet (zhang2023adding), others incorporate explicit constraints such as segmentation masks (wang2024gaussianeditor) or cross-attention maps (zhuang2023dreameditor). However, these methods struggle with 3D object insertion, which demands reasoning about semantically appropriate yet physically unoccupied regions for placement. In this work, we primarily focus on the task of object insertion in 3D scenes.
@@
Object Insertion in 3D Scene. In contrast to modifying existing scene content, object insertion remains underexplored. MVInpainter (caomvinpainter) leverages segmentation to identify support regions like table surfaces, but struggles with fine-grained insertions on objects or humans. GaussianEditor (chen2024gaussianeditor) and InseRF (shahbazi2024inserf) insert objects using 3D reconstruct models, yet still require user-provided 2D masks and suffer from depth-related localization issues. FocalDreamer (li2024focaldreamer) attaches parts to base shapes but depends on user-specified 3D parameters (e.g., rotation, translation, scale), and lacks generalization to complex scenes. Other methods (zhuang2024tip; shum2024language) guide object generation via diffusion using 3D bounding boxes, which imposes a burden on users. In this work, we aim for unsupervised and broadly applicable 3D object insertion, removing the need for manual annotations or spatial priors.
@@
Large Language Models in 3D Generation and Editing. LLMs, like GPT (achiam2023gpt) and Llama (grattafiori2024llama) series, have exhibited outstanding efficacy in many text-related tasks. zhou2024gala3d and zhou2025layoutyourd utilize LLMs to provide coarse compositional spatial information from textual descriptions to construct the 3D scene. The multi-modal variants of LLMs (deitke2024molmo; achiam2023gpt) incorporate images and are additionally trained on image-text pairs, showing impressive results for visual captioning and vision question-answering (VQA). Notably, Molmo (deitke2024molmo) can perform pixel-level localization mainly because it was trained with richly annotated image data. This capability is crucial for robust spatial grounding in vision-language tasks. GG-Editor (xu2024gg) first exploits GPT-4V (achiam2023gpt) to better understand both the textual and 3D visual inputs and then infer reasonable local regions for 3D editing. However, it primarily targets object editing and its preliminary use of MLLM struggles to ensure spatial precision. In this work, we leverage the text reasoning capabilities and spatial relationship understanding of GPT-4 (achiam2023gpt) and Molmo (deitke2024molmo), supplemented by a basic detection model (xiao2024florence), to eliminate the reliance on manually provided priors in object insertion.
@@ 3.2. MLLM-based Object Insertion Parser
A key challenge in unsupervised object insertion is converting high-level user intent into structured, fine-grained guidance. To address this, we introduce an MLLM-based Object Insertion Parser (MLLM-OIP) that utilizes the MLLM’s spatial understanding capability to parse the instruction $\mathcal{T}$ into semantically prompts, providing essential guidance for the subsequent object insertion. Specifically, we provide a prompt template $\mathcal{T}_{\textit{parser}}$ and a sampled scene image $\mathcal{I}_{\textit{S}}$ as input to the multimodal LLM $\mathcal{M}_{\textit{MLLM}}$ (achiam2023gpt) to obtain structured outputs. The prompts generation process is formalized as :
@@ 3.3. Initialization via Large Models
Object from Prompts. To avoid 3D inconsistency, we first use a text-to-image (T2I) (rombach2022high) model to synthesize a Text-generated image $\mathcal{I}_{\textit{O}}$ of the object from the object description prompt $\mathcal{T}_{\textit{O}}$. The synthesized image is then used to recover the 3D geometry $\mathcal{G}_{O}$ via LGM (tang2024lgm), a single-view reconstruction model that achieve a trade-off between reconstruction quality and efficiency. Other lightweight 3D reconstruction methods (long2024wonder3d; hong2023lrm) can also be adopted. In addition, $\mathcal{I}_{\textit{O}}$ can be directly specified by the user, allowing for more precise control over the object’s appearance.
Scene’s Attachment Region Detection. Intuitively, an object’s placement is influenced by the attachment region of the scene and the degrees of freedom within that region. Driven by this, we extract an attachment region $\mathcal{G}_{\textit{AR}}$ and its associated degrees of freedom $\textit{DoF}_{\textit{AR}}$($\textit{s}_{\textit{AR}}$, $\textit{r}_{\textit{AR}}$, $\textit{t}_{\textit{AR}}$) from the 3D scene based on the Attachment Region Prompt $\mathcal{T}_{\textit{AR}}$. This region serves as a crucial spatial reference, guiding initializing the inserted object $\textit{DoF}_{\textit{O}}$. Specifically, we employ an open-vocabulary detection model Florence2 (xiao2024florence) to localize 2D bounding boxes across smapled views from the scene with camera poses $\mathcal{C}_{\textit{cam}}$, guided by $\mathcal{T}_{\textit{AR}}$. For each view, the detected box is converted into a binary mask $\mathcal{I}_{\textit{BAR}}$, representing the candidate attachment region. The 3D attachment area is parameterized by the degrees of freedom of a initial 3D bounding box $\mathcal{B}_{\textit{init}}$. We optimize the attachment by computing cross-entropy between the projected transformed bounding box and the detected attachment region mask $\mathcal{B}_{\text{AR}}$ across all camera views. Thus, $\textit{DoF}_{\textit{AR}}$ is calculated as follows:
@@
To initialize the $\textit{t}_{\textit{O}}$, we use strong pixel-level semantic spatial localization capability of Molmo (deitke2024molmo) to predict a set of 2D object centers $\{c_{O}^{(v)}\}_{v\in\mathcal{V}}$ across multiple scene views with the Local Target $\mathcal{T}_{\textit{LT}}$, using prompt like “Point the position to add $\mathcal{T}_{\textit{LT}}$¿”. Let $\hat{\mathcal{G}}_{O}^{(t)}$ denote the object geometry after applying the transformation $(\textit{s}_{\textit{O}}$, $\textit{r}_{\textit{O}}$, $t$ ), where $t$ is a optimized parameter. For each view $v$, we project the transformed object and compute the 2D centroid of its projection. The $\textit{t}_{\textit{O}}$ is obtained by optimizing $t$ to minimize the discrepancy between the projected and the predicted centroids:
@@
Here, $\pi_{v}(\cdot)$ is the camera projection function for view $v$, and Centroid ($\cdot$) computes the 2D projected center of the object. To ensure physical plausibility during object insertion, we introduce the collision loss (zhou2025layoutyourd) $\mathcal{L}_{\text{coll}}$, which penalizes interpenetration between the object centriod $O_{c}$ and the scene attachment region $\mathcal{G}_{\text{AR}}$.
@@ 3.4. Hierarchical Spatial Aware Refinement
The initial $\textit{DoF}_{\textit{O}}$ from $\mathcal{M}_{\textit{MLLM}}$ often lack spatial accuracy, hindering seamless scene integration. Base on that, we then optimize the $\textit{DoF}_{\textit{O}}$ using SSDS Loss (chen2024comboverse), refining the object’s placement in the scene. The loss is defined as:
@@
Attention-based Localization. We found that due to the bias in the T2I model’s training data, SSDS Loss exhibits limitations when handling rare spatial relationships. These inherent limitations restrict the effectiveness of prompt instructions. To enable stronger spatial conditioning, we adopt attention-based localization loss (zhuang2024tip), enforcing tighter regional constraints as follows: